# Analyzing the Generalization and Reliability of Steering Vectors

### v1

@@ Abstract
Steering vectors (SVs) are a new approach to efficiently adjust language model behaviour at inference time by intervening on intermediate model activations. They have shown promise in terms of improving both capabilities and model alignment. However, the reliability and generalisation properties of this approach are unknown. In this work, we rigorously investigate these properties, and show that steering vectors have substantial limitations both in- and out-of-distribution. In-distribution, steerability is highly variable across different inputs. Depending on the concept, spurious biases can substantially contribute to how effective steering is for each input, presenting a challenge for the widespread use of steering vectors. Out-of-distribution, while steering vectors often generalise well, for several concepts they are brittle to reasonable changes in the prompt, resulting in them failing to generalise well. Overall, our findings show that while steering can work well in the right circumstances, there remain many technical difficulties of applying steering vectors to guide models’ behaviour at scale.
Machine Learning, ICML
@@ 1 Introduction
Steering Vectors (SVs) (Rimsky et al., 2023; Turner et al., 2023; Zou et al., 2023; Liu et al., 2023) have been recently proposed as a technique for guiding language model behaviour at inference time. Existing work has shown promising results in using these SVs to detect and guide models towards high-level traits such as honesty (Zou et al., 2023), sycophancy (Rimsky et al., 2023), and positive sentiment (Tigges et al., 2023). They have also been shown to be useful for improving model capabilities (Wu et al., 2024; Liu et al., 2023; van der Weij et al., 2024) and red-teaming (Rimsky, 2023). SVs are of interest as they enjoy a number of practical benefits over other model adjustment techniques that require adding more information into the context window (Brown et al., 2020; Wei et al., 2022, In-Context Learning), or performing training to adjust model parameters (fine-tuning). Recent work shows that steering vectors can be learned in an unsupervised way (Mack and Turner, 2024), thus removing another obstacle for their use. It may even be possible for different steering vectors to be used in combination for multiple behaviours (van der Weij et al., 2024; Wang et al., 2023b). It would thus be very important and useful in practice if steering vectors were truly effective.
However, existing work has mostly evaluated SVs in-distribution, and looked at aggregate behaviour. It is unknown how reliable the change in behaviour caused by SVs is, and how well SVs generalise to different system or user prompts. In this paper, we extensively evaluate the in-distribution reliability and out-of-distribution generalisation of SVs, extending analysis in Rimsky et al. (2023) to a much broader variety of behaviours from the Model Written Evals (MWE) datasets by Perez et al. (2022). In addition, we consider targeted distribution shifts in the form of inserting prompts via the user message or system message. This setting mimics the practically important setting where we will need to apply SVs to different system and user prompts, and where we would require SVs to generalise well to be robustly useful.
Our first key result is that for many behaviours studied, steering is unreliable (Figures 1 and 5). For all behaviours evaluated, steerability takes on a large range of values across different inputs, including negative values, where SVs produce the opposite of the desired behaviour. Previous work (Rimsky et al., 2023; Turner et al., 2023; Zou et al., 2023; Liu et al., 2023) does not study this variance, which potentially leads to over-optimistic claims on performance due to a lack of error bars. In explaining this variance, we demonstrate a novel type of bias, steerability bias, in which models are easier to steer towards outputs with a certain property (i.e. answer position or token choice). The lack of steerability and high variance in steering performance demonstrates that in many cases, a steering vector extracted may not correspond to the intended concept, and applying steering vectors may only be effective in the presence of spurious factors associated with the prompt template or other potential biases.
@@ 2.1 Steering Vectors
Steering Vectors (SVs, also known as activation engineering) and related ideas were introduced by Turner et al. (2023); Zou et al. (2023); Liu et al. (2023). SVs can be seen as an inference-time intervention (Li et al., 2023) technique in the representation engineering (Zou et al., 2023) toolkit, which is an umbrella term for the broad approach of improving the transparency and controllability of neural networks by examining and intervening on population-level representations and activations of the network. (Rajendran et al., 2024). Rimsky et al. (2023) recently introduced Contrastive Activation Addition (CAA), a specific technique for extracting and applying SVs which we use in this work, due to its simplicity, effectiveness and popularity in the community. Rimsky et al. (2023) demonstrate the effectiveness of CAA in-distribution on several AI alignment-relevant behaviours, while we test on a much broader range of behaviours, investigate the reliability of the steering intervention, and examine out-of-distribution generalisation of SVs. We describe the CAA method in more detail in Section 3.
@@
Compared to fine-tuning (Ziegler et al., 2019; Rafailov et al., 2023; Mishra et al., 2022), steering vectors don’t involve changing model parameters, hence potentially avoiding catastrophic forgetting (Biderman et al., 2024; Luo et al., 2023). Compared to in-context learning (Brown et al., 2020; Wei et al., 2022, ICL), steering does not require adding tokens to the prompt, saving inference cost and enabling it to scale beyond the length of the context window. Furthermore, Zou et al. (2023) show that steering interventions are robust to adversarial attacks capable of breaking prompt-based and fine-tuning-based alignment methods (Perez et al., 2022; Wei et al., 2023; Jiang et al., 2024).
2.2 Steering Vectors and the Linear Representation Hypothesis
The effectiveness of steering vectors in- and out-of-distribution has implications for the linear representation hypothesis (LRH) (Park et al., 2023). A key prediction of the LRH is that each atomic feature is associated with a single global direction in activation space, and that intervening by adding or subtracting this direction can influence the model’s understanding and / or behaviour. Previous work that validates the LRH mostly considers the in-distribution (ID) setting (Mallen and Belrose, 2023; Burns et al., 2022; Marks and Tegmark, 2023; Nanda et al., 2023; Li et al., 2022). However, this is only evidence of local linearity, which is satisfied by all continuous functions within a sufficiently small neighbourhood. The LRH in fact makes a stronger claim: that representations are globally linear. For SVs to generalise well OOD, this stronger claim has to be true — although it may not be sufficient, and the reverse implication doesn’t necessarily hold, as the concepts that are linearly represented might not be human-interpretable or extractable with SV approaches.
Therefore, our analysis can be seen as extending existing validations of the LRH to the more challenging out-of-distribution (OOD) setting. Crucially, our proposed experimental protocol can differentiate the LRH from competing frameworks which allow for local, but not global, linearity (Elhage et al., 2022; Black et al., 2022). While we primarily focus on the practical usefulness of SVs in this paper, we want to highlight the additional value of our work from the perspective of the LRH.
2.3 Evaluating Generalisation Behaviour of Model Adjustment Procedures
Several existing works evaluate the generalisation properties of model adjustment techniques. Hupkes et al. introduce a taxonomy of work investigating generalisation in NLP. In the language of their taxonomy, the investigation in Section 6 has a practical motivation and uses a generated covariate shift for testing generalisation across domains at the fine-tuning (or in our case SV-training) stage. Kirk et al. (2023) investigate the generalisation properties of different fine-tuning approaches in summarisation and instruction-following settings. Clymer et al. (2023) investigate generalisation of different methods for training reward models across a range of generated natural shifts, and find that methods based on similar ideas to SVs often generalise reasonably well. In contrast, our work investigates the generalisation properties of SVs specifically on a wide range of datasets that capture desirable properties we would want to steer models towards or away from, and investigates generalising over different prompts, rather than different inputs.
@@ 3 Preliminaries
Rimsky et al. (2023) propose Contrastive Activation Addition (CAA) to extract and apply steering vectors on datasets. We follow this protocol in our experiments, and so we summarise the main steps here.
@@
We note that other aggregation methods have been proposed, but literature does not suggest these perform better than mean-difference. We discuss alternatives in Section C.4.
@@ 4.1 Datasets and Prompts
We focus on the Model-Written Evaluations (MWE) datasets (Perez et al., 2022), a large dataset consisting of prompts from over 100 distinct categories designed to evaluate many specific aspects of models’ behaviour. Each category contain 1000 samples generated by an LLM, covering a variety of persona and behaviors. For each of these datasets, we construct a 40-10-50 train-val-test split. We also include TruthfulQA (Lin et al., 2022) and the sycophancy dataset (Perez et al., 2022), as they were used in CAA (Rimsky et al., 2023). The validation split is used for hyperparameter selection; we discuss this in Section 4.3. We randomly choose three persona datasets from each MWE persona dataset category, while keeping the sycophancy, TruthfulQA, and AI risk datasets used in CAA for a total of 40 datasets.
@@ 4.2 Metrics
To measure the effectiveness of steering, we need a metric of the model’s propensity to exhibit a behavioural trait (e.g. sycophancy (Rimsky et al., 2023), truthfulness (Lin et al., 2022), helpfulness (Bai et al., 2022)). Given a propensity metric, we then define propensity curves and steerability as summary metrics of the steering vector’s effectiveness.
@@
Rimsky et al. (2023) instead uses the normalised probability of the positive answer, which is the same except for a softmax applied to the logits. We note that normalised probabilities are a monotonic function of the logit difference, so propensity is order-invariant between these two methods. However, logit-difference is likely to be more linear with respect to the model’s intermediate activations (as it doesn’t include a softmax), facilitating downstream analysis.
Figure 2: Example propensity curve (dots) and linear steerability fit. The slope of the linear fit is the steerability.
@@
To get a sense of how well steering works as a function of the multiplier $\lambda$, we compute $m_{LD}$ for various values of $\lambda\in\Lambda=\{-1.5,-1.0,-0.5,0.0,0.5,1.0,1.5\}$. We refer to this as a propensity curve, which was proposed by Rimsky et al. (2023). If steering works well, we expect the trend to be monotonic and increasing.
@@ 4.3 Steering Vector Extraction
Following previous work, we focus on steering instruction-tuned models. We include Llama-2-7b-Chat (Touvron et al., 2023) as it was used in previous work. In order to draw conclusions that generalise beyond a single model, we also consider Qwen-1.5-14b-Chat (Bai et al., 2023), which differs in many aspects, including architecture, parameter count, and training data distributions.
@@
The choice of which layer to steer at is an important hyperparameter. Loosely, we expect that each layer captures a different level of abstraction in the model’s internal computation (Elhage et al., 2021), and steering will work best if we choose the layer that best matches the target concept’s level of abstraction. In order to determine the optimal layer, we sweep over all layers using the validation split. In line with Rimsky et al. (2023), we find that the optimal choice of layer is remarkably consistent across many datasets. Thus, we fix layer 13 for Llama and layer 21 for Qwen for all subsequent experiments. Layer response curves used in selecting the optimal layer are presented in Section C.7.
@@ 5 Evaluating Steering Vector Reliability
This bias is different from the standard position or token bias exhibited by LLMs (Zheng et al., 2023; Wang et al., 2023a), as it is a steerability bias: the model is more steerable towards the positive answer when it a particular position or token compared to the other position or token. The preferred token or position is different for each dataset; for example corrigible-neutral-HHH has a B-steerability bias, whereas self-aware-lm has an A-steerability bias (see Figure 3). This is problematic, as it is not fixable by simple dataset debiasing (which was already performed) or logit calibration adjustments (Zhou et al., 2023) (as they effect propensity, not the change in propensity, i.e. steerability). Further, it implies that there may be other steerability biases present in models, determining when they are more or less steerable towards specific answers or behaviours. Indeed, there is still a high degree of unexplained variance present in many datasets in Figure 4.
@@ 6 Steering Out-of-Distribution
We compare aggregate in-distribution and out-of-distribution steerability between Llama and Qwen in Figure 7. We find that both ID and OOD steerabilities are highly correlated across models, despite them having different sizes, architectures, and training procedures. The consistency between different model architectures indicates that the effectiveness of a steering vector is mostly a property of the dataset used to extract the dataset, as opposed to the model used. This may also be evidence that different models converge to similar ontologies (Huh et al., 2024; Gurnee et al., 2024).
@@
While steerability and SV generalisation is mostly a dataset-level property, there is still variation in generalisation performance that is not captured by dataset; for example, SVs generalise better over some shifts than others for the same dataset. In Figure 8 show that the similarity in the propensity of the model in two prompt settings is correlated with the relative steerability (Equation 2), a measure of generalisation. In other words, if the model behaves similar in two prompt settings, then SVs will transfer better between those two settings than if the model behaves differently in the two settings. We show a similar result but for SV cosine similarity in Section D.4.
@@ 7 Discussion and Conclusion
Our work is the first to report and analyse the variance in steerability at a per-example level, and in doing so reveal a major limitation in SV reliability. In Section 5, we demonstrated that the reliability of SVs’s effects on model behaviour are highly dataset-dependent, with some concepts being unsteerable and some SVs producing the opposite behaviour to what is desired. We found that this unreliability is often driven by token- and position-steerability bias, a new type of bias we discovered that is distinct from standard token and position biases in LLMs. Although these are very simple biases which can be easily understood, simple interventions in data preprocessing fail to address the problem, and there are likely to be other steerability biases that will affect the effectiveness and reliability of SVs. In Section D.3 we show that this variance is partially a dataset property rather than a model property, implying that future work investigating what causes these biases should at least partially focus on the dataset, as well as analysing whether other techniques for extracting and applying SVs can mitigate these biases.
In Section 6 we evaluated the generalisation properties of SVs, finding that while they often generalise reasonably well (conditioned on their in-distribution performance being good), generalisation is not always perfect. We find that SV generalisation is mostly a property of the dataset, and is correlated by the similarity in un-steered propensity of the model in the source and target setting. This correlation is problematic, as often we would want to apply steering vectors to guide model behaviour towards something it does not normally do, but in these scenarios SVs tend to generalise less well. Investigating methods to improve SV generalisation, and investigating scenarios where they generalise better or worse is important future work. Furthermore, concurrent work suggests that steering for one behaviour results in trade-offs with other behaviours; future work could elucidate this by extending our analysis to cross-concept steering.
@@
8 Author Contribution Statement
Daniel Tan and David Chanin were responsible for building out the codebase and running experiments. Daniel Tan and Robert Kirk were the main writers of the paper. Robert Kirk was the main technical supervisor, providing frequent and intensive supervision. Aengus Lynch and Adria Garriga-Alonso helped with editing and provided useful feedback / advice. Brooks Paige and Dimitrios Kanoulas provided useful feedback.
9 Acknowledgements
We thank Nina Rimsky, Teun van der Weij, and Felix Hofstatter for useful feedback and advice on the paper. Daniel is supported by a scholarship from the Agency for Science, Technology, and Research (Singapore).
@@ Appendix A Hardware Requirements
Appendix B Limitations

### v4

# Analysing the Generalisation and Reliability of Steering Vectors

@@ Abstract
Steering vectors (SVs) are a new approach to efficiently adjust language model behaviour at inference time by intervening on intermediate model activations. They have shown promise in terms of improving both capabilities and model alignment. However, the reliability and generalisation properties of this approach are unknown. In this work, we rigorously investigate these properties, and show that steering vectors have substantial limitations both in- and out-of-distribution. In-distribution, steerability is highly variable across different inputs. Depending on the concept, spurious biases can substantially contribute to how effective steering is for each input, presenting a challenge for the widespread use of steering vectors. We additionally show steerability is also mostly a property of the dataset rather than the model by measuring steerability across multiple models. Out-of-distribution, while steering vectors often generalise well, for several concepts they are brittle to reasonable changes in the prompt, resulting in them failing to generalise well. Similarity in behaviour between distributions somewhat predicts generalisation performance, but there is more work needed to understand when and why steering vectors generalise correctly. Overall, our findings show that while steering can work well in the right circumstances, there remain many technical difficulties of applying steering vectors to robustly guide models’ behaviour at scale.
@@ 1 Introduction
Steering Vectors (SVs) [30, 33, 43, 18] have been recently proposed as a technique for guiding language model behaviour at inference time. Existing work has shown promising results in using these SVs to detect and guide models towards high-level traits such as honesty [43], sycophancy [30], and positive sentiment [31]. They have also been shown to be useful for improving model capabilities [39, 18, 34] and red-teaming [29]. SVs are of interest as they enjoy a number of practical benefits over other model adjustment techniques that require adding more information into the context window [5, 38, In-Context Learning], or performing training to adjust model parameters (fine-tuning). Recent work shows that steering vectors can be learned in an unsupervised way [20], thus removing another obstacle for their use. It may even be possible for different steering vectors to be used in combination for multiple behaviours [34, 36]. It would thus be very important and useful in practice if steering vectors were truly effective.
However, existing work has mostly evaluated SVs in-distribution, and looked at aggregate behaviour. It is unknown how reliable the change in behaviour caused by SVs is, and how well SVs generalise to different system or user prompts. In this paper, we extensively evaluate the in-distribution reliability and out-of-distribution generalisation of SVs, extending analysis in Rimsky et al. [30] to a much broader variety of behaviours from the Model Written Evals (MWE) datasets by Perez et al. [26]. In addition, we consider targeted distribution shifts in the form of inserting prompts via the user message or system message. This setting mimics the practically important setting where we will need to apply SVs to different system and user prompts, and where we would require SVs to generalise well to be robustly useful.
Our first key result is that for many behaviours studied, steering is unreliable (Figures 1 and 5). For all behaviours evaluated, steerability takes on a large range of values across different inputs, including negative values, where SVs produce the opposite of the desired behaviour. Previous work [30, 33, 43, 18] does not study this variance, which potentially leads to over-optimistic claims on performance due to a lack of error bars. In explaining this variance, we demonstrate a novel type of bias, steerability bias, in which models are easier to steer towards outputs with a certain property (i.e. answer position or token choice). The lack of steerability and high variance in steering performance demonstrates that in many cases, a steering vector extracted may not correspond to the intended concept, and applying steering vectors may only be effective in the presence of spurious factors associated with the prompt template or other potential biases.
@@ 2 Related Work
Steering Vectors (SVs, also known as activation engineering) and related ideas were introduced by Turner et al. [33], Zou et al. [43], Liu et al. [18]. SVs can be seen as an inference-time intervention [16] technique in the representation engineering [43] toolkit, which is an umbrella term for the broad approach of improving the transparency and controllability of neural networks by examining and intervening on population-level representations and activations of the network. [28]. Rimsky et al. [30] recently introduced Contrastive Activation Addition (CAA), a specific technique for extracting and applying SVs which we use in this work, due to its simplicity, effectiveness and popularity in the community. Rimsky et al. [30] demonstrate the effectiveness of CAA in-distribution on several AI alignment-relevant behaviours, while we test on a much broader range of behaviours, investigate the reliability of the steering intervention, and examine out-of-distribution generalisation of SVs. We describe the CAA method in more detail in Section 3.
@@
Compared to fine-tuning [42, 27, 23], steering vectors don’t involve changing model parameters, hence potentially avoiding catastrophic forgetting [3, 19]. Compared to in-context learning [5, 38, ICL], steering does not require adding tokens to the prompt, saving inference cost and enabling it to scale beyond the length of the context window. Furthermore, Zou et al. [43] show that steering interventions are robust to adversarial attacks capable of breaking prompt-based and fine-tuning-based alignment methods [26, 37, 13]. An extended related work section including discussing the relationship between SVs and the Linear Representation Hypothesis [25, LRH] and other works that evaluate the generalisation behaviour of model adjustment methods can be found in Appendix B.
@@ 3 Preliminaries
Rimsky et al. [30] propose Contrastive Activation Addition (CAA) to extract and apply steering vectors on datasets. We follow this protocol in our experiments, and so we summarise the main steps here.
@@
We note that other aggregation methods have been proposed, but literature does not suggest these perform better than mean-difference. We discuss alternatives in Section D.4.
@@ 4.1 Datasets and Prompts
We focus on the Model-Written Evaluations (MWE) datasets [26], a large dataset consisting of prompts from over 100 distinct categories designed to evaluate many specific aspects of models’ behaviour. Each category contain 1000 samples generated by an LLM, covering a variety of persona and behaviors. For each of these datasets, we construct a 40-10-50 train-val-test split. We also include TruthfulQA [17] and the sycophancy dataset [26], as they were used in CAA [30]. The validation split is used for hyperparameter selection; we discuss this in Section 4.3. We randomly choose three persona datasets from each MWE persona dataset category, while keeping the sycophancy, TruthfulQA, and AI risk datasets used in CAA for a total of 40 datasets.
@@ 4.2 Metrics
To measure the effectiveness of steering, we need a metric of the model’s propensity to exhibit a behavioural trait (e.g. sycophancy [30], truthfulness [17], helpfulness [2]). Given a propensity metric, we then define propensity curves and steerability as summary metrics of the steering vector’s effectiveness.
@@
Rimsky et al. [30] instead uses the normalised probability of the positive answer, which is the same except for a softmax applied to the logits. We note that normalised probabilities are a monotonic function of the logit difference, so propensity is order-invariant between these two methods. However, logit-difference is likely to be more linear with respect to the model’s intermediate activations (as it doesn’t include a softmax), facilitating downstream analysis.
@@
Figure 2: Example propensity curve and steerability fit for high steerability (left), and low (right).
@@
To get a sense of how well steering works as a function of the multiplier $\lambda$, we compute $m_{LD}$ for various values of $\lambda\in\Lambda=\{-1.5,-1.0,-0.5,0.0,0.5,1.0,1.5\}$. We refer to this as a propensity curve, which was proposed by [30]. If steering works well, we expect the trend to be monotonic and increasing with high slope.
@@ 4.3 Steering Vector Extraction
Following previous work, we focus on steering instruction-tuned models. We include Llama-2-7b-Chat [32] as it was used in previous work. In order to draw conclusions that generalise beyond a single model, we also consider Qwen-1.5-14b-Chat [1], which differs in many aspects, including architecture, parameter count, and training data distributions.
@@
The choice of which layer to steer at is an important hyperparameter. Loosely, we expect that each layer captures a different level of abstraction in the model’s internal computation [8], and steering will work best if we choose the layer that best matches the target concept’s level of abstraction. In order to determine the optimal layer, we sweep over all layers using the validation split. In line with Rimsky et al. [30], we find that the optimal choice of layer is remarkably consistent across many datasets. Thus, we fix layer 13 for Llama and layer 21 for Qwen for all subsequent experiments. Layer response curves used in selecting the optimal layer are presented in Section D.7.
@@ 5 Evaluating Steering Vector Reliability
This bias is different from the standard position or token bias exhibited by LLMs [40, 35], as it is a steerability bias: the model is more steerable towards the positive answer when it a particular position or token compared to the other position or token. The preferred token or position is different for each dataset; for example corrigible-neutral-HHH has a B-steerability bias, whereas self-aware-lm has an A-steerability bias (see Figure 14). This is problematic, as it is not fixable by simple dataset debiasing (which was already performed) or logit calibration adjustments [41] (as they effect propensity, not the change in propensity, i.e. steerability). Further, it implies that there may be other steerability biases present in models, determining when they are more or less steerable towards specific answers or behaviours. Indeed, there is still a high degree of unexplained variance present in many datasets in Figure 4.
@@ 6 Steering Out-of-Distribution
We compare aggregate in-distribution and out-of-distribution steerability between Llama and Qwen in Figure 7. We find that both ID and OOD steerabilities are highly correlated across models, despite them having different sizes, architectures, and training procedures. The consistency between different model architectures indicates that the effectiveness of a steering vector is mostly a property of the dataset used to extract the dataset, as opposed to the model used. This may also be evidence that different models converge to similar ontologies [11, 10].
@@
While steerability and SV generalisation is mostly a dataset-level property, there is still variation in generalisation performance that is not captured by dataset; for example, SVs generalise better over some shifts than others for the same dataset. In Figure 8 show that the similarity in the propensity of the model in two prompt settings is correlated with the relative steerability (Equation 2), a measure of generalisation. In other words, if the model behaves similar in two prompt settings, then SVs will transfer better between those two settings than if the model behaves differently in the two settings. We show a similar result but for SV cosine similarity in Section E.3.
@@ 7 Discussion and Conclusion
Our work is the first to report and analyse the variance in steerability at a per-example level, and in doing so reveal a major limitation in SV reliability. In Section 5, we demonstrated SVs’s effects on model behaviour are often unreliable, with some concepts being unsteerable and some SVs producing the opposite behaviour to what is desired. We found that this unreliability is often driven by token- and position-steerability bias, a new type of bias we discovered that is distinct from standard token and position biases in LLMs. Although these are very simple biases which can be easily understood, simple interventions in data preprocessing fail to address the problem, and there are likely to be other steerability biases that will affect the effectiveness and reliability of SVs. In Section E.2 we show that this variance is partially a dataset property rather than a model property, implying that future work investigating what causes these biases should at least partially focus on the dataset, as well as analysing whether other techniques for extracting and applying SVs can mitigate these biases.
In Section 6 we evaluated the generalisation properties of SVs, finding that while they often generalise reasonably well (conditioned on their in-distribution performance being good), generalisation is not always perfect. We find that SV generalisation is mostly a property of the dutataset, and is correlated by the similarity in un-steered propensity of the model in the source and target setting. This correlation is problematic, as often we would want to apply steering vectors to guide model behaviour towards something it does not normally do, but in these scenarios SVs tend to generalise less well. Investigating methods to improve SV generalisation, and investigating scenarios where they generalise better or worse is important future work.
@@ Appendix A Hardware Requirements
Appendix B Extended Related Work
B.1 Steering Vectors and the Linear Representation Hypothesis
The effectiveness of steering vectors in- and out-of-distribution has implications for the linear representation hypothesis (LRH) [25]. A key prediction of the LRH is that each atomic feature is associated with a single global direction in activation space, and that intervening by adding or subtracting this direction can influence the model’s understanding and / or behaviour. Previous work that validates the LRH mostly considers the in-distribution (ID) setting [21, 6, 22, 24, 15]. However, this is only evidence of local linearity, which is satisfied by all continuous functions within a sufficiently small neighbourhood. The LRH in fact makes a stronger claim: that representations are globally linear. For SVs to generalise well OOD, this stronger claim has to be true — although it may not be sufficient, and the reverse implication doesn’t necessarily hold, as the concepts that are linearly represented might not be human-interpretable or extractable with SV approaches.
Therefore, our analysis can be seen as extending existing validations of the LRH to the more challenging out-of-distribution (OOD) setting. Crucially, our proposed experimental protocol can differentiate the LRH from competing frameworks which allow for local, but not global, linearity [9, 4]. While we primarily focus on the practical usefulness of SVs in this paper, we want to highlight the additional value of our work from the perspective of the LRH.
B.2 Evaluating Generalisation Behaviour of Model Adjustment Procedures
Several existing works evaluate the generalisation properties of model adjustment techniques. Hupkes et al. [12] introduce a taxonomy of work investigating generalisation in NLP. In the language of their taxonomy, the investigation in Section 6 has a practical motivation and uses a generated covariate shift for testing generalisation across domains at the fine-tuning (or in our case SV-training) stage. Kirk et al. [14] investigate the generalisation properties of different fine-tuning approaches in summarisation and instruction-following settings. Clymer et al. [7] investigate generalisation of different methods for training reward models across a range of generated natural shifts, and find that methods based on similar ideas to SVs often generalise reasonably well. In contrast, our work investigates the generalisation properties of SVs specifically on a wide range of datasets that capture desirable properties we would want to steer models towards or away from, and investigates generalising over different prompts, rather than different inputs.
Appendix C Limitations